Inductance of moving iron instrument is given by L=(12+6ø-ø2)µH Where ø is the Deflection in radians from zero position. The spring constant is 12*10-6 Nm/Rad. calculate the Deflection for current of 8 amp.


1
Expert's answer
2021-12-21T02:04:06-0500

For the given current, the deflection is


"\theta=\frac12 \frac{I^2}{k}\frac{\text dL}{\text d\theta},\\\space\\
\theta=\frac12 \frac{I^2}{k}\frac{\text d(12+6\theta-\theta^2)}{\text d\theta},\\\space\\
\theta=\frac12 \frac{I^2}{k}(6-2\theta),\\\space\\
\theta=3\text{ rad}."